what is the smallest positive number for t?
sin(t)+cos(2t)=0

Answer by lwsshak3(11628)

what is the smallest positive number for t?
sin(t)+cos(2t)=0
sint+cos^2t-sin^2t=0
sint+1-sin^2t-sin^2t=0
2sin^2t-sint-1=0
(2sint+1)(sint-1)=0
sint-1=0
sint=1
t=π/2
or
2sint+1=0
sint=-1/2
t=7π/4
..
smallest positive number for t=π/2